a1413

Hb (21,632 GB on HOB+0127)

  • Experiment started OK
  • Missed scans 262-0202 through 262-0224 due to system check

Ke (23,354 GB on HOB+0125)

  • Experiment started OK

Yg (11,937 GB on HOB+1007)

  • Experiment started OK
  • Missed scans 262-0329b through 262-0354 due to system check.
  • Module swapped over by itself at 20:26UT, now recording to fresh 16TB module (on this scan - 262-2026 9102 20:26:37 20:28:17 4149.2)